OpenStudy (anonymous):

Log subscript 7(8x+20)=Log subscript 7(x+6) . Solve for x

OpenStudy (hunus):

If\[\Large \log_{7}(8x+20)=\log_{7}(x+6)\]then\[\Large 8x+20=x+6\]

OpenStudy (whpalmer4):

Use the property of logarithms that \(b^{\log_b{a}} = a\)

OpenStudy (anonymous):

So now how do you solve ?

OpenStudy (hunus):

Do you know how to solve 8x + 20 = x + 6 ?

OpenStudy (anonymous):

got x=-2

OpenStudy (hunus):

That's your answer

OpenStudy (whpalmer4):

That's correct!

OpenStudy (anonymous):

log subscript4(4x) + log subscript 4(x)= Log subscript 4(64)

OpenStudy (anonymous):

solve for x

OpenStudy (whpalmer4):

\[\log{a}+\log{b} = \log{(a*b)}\]

OpenStudy (whpalmer4):

So you can rewrite as \(\log_4{(4x*x)} = \log_4(64)\) or \(\log_4{(4x^2)} = \log_4(64)\) From there you should be able to solve as in the previous problem.

OpenStudy (anonymous):

x=4

OpenStudy (whpalmer4):

Let's check: \[\log_4{4(4)} + \log_4{4} = \log_4{64}\]\[\log_4 16 + \log_4 4 = \log_4 64\]\[2+1=3\checkmark\]